OpenStudy (anonymous):

Quantity x minus 1 all over 9 = 2 for x.

OpenStudy (compassionate):

\[\frac{ x - 1 }{ 9 } = 2\] Start by performing the opposite action on 9. That is, multiply both sides by 9. x - 1 = 18. Now x = 18 + 1